# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 4.2.2. Neck-stretching
    The period of $\gamma_{p}$ is $1/f(p_{i})$ (which is approximately $1$ by choice of $f$), and hence $\gamma_{p}$ has longer period than $\gamma_{q}$ if and only if $f(p)<f(q)$.
@@
    The set of all Reeb orbits with period at most $2$ is just $\{\,\gamma_{p}\;|\;p\in\mathrm{Crit}(f)\,\}$.
Now, $M\#\partial(V\times\mathbb{D}^{2})$ is obtained by contact surgery of index 00 on the disjoint union $M\sqcup\partial(V\times\mathbb{D}^{2})$, or, in other words, it is the contact boundary of a Weinstein $1$-handle attachment on the disjoint union of the Liouville filling $F$ of $M$ with $V\times\mathbb{D}^{2}$. The part of surgery (or handle attachment) in the $\partial(V\times\mathbb{D}^{2})$ component can be more precisely done inside the subset $V\times S^{1}\subset\partial(V\times\mathbb{D}^{2})$, away from every critical point of $f$ (times the $S^{1}$ factor), and more precisely in the region $f^{-1}((a,b))$, where $a$ is the minimum of $f$ and $b$ is the image via $f$ of its index $1$ critical points (recall $f$ is self indexing, so all index 00 and index $1$ critical points have same image). We use here in particular the special contact form, that we denote $\lambda$, on the connected sum given by [Laz20a, Theorem 3.15] and [Yau04] as done for Case 1. More precisely, the action of the new simple closed Reeb orbit $\gamma_{h}$ on the belt sphere of the resulting connect sum depends on the thickness of the Weinstein handle realizing this contact surgery, and we fix once and for all such thickness to be very small (which results in the action of $\gamma_{h}$ to also be very small). According to [Laz20a, Theorem 3.15] and [Yau04], one can in fact arrange that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period $<2$ are just the $\gamma_{p}$’s and (some of) the multiple covers of the simple Reeb orbit $\gamma_{h}$ in the belt sphere of the handle.
In our setting, we moreover claim that we can attach the $1$-handle to the $V\times S^{1}$ region in such a way that the contact form $\lambda$ satisfies the following property: whenever $p_{i}$ is not the minimum point of $f$,
$\frac{1}{f(p_{i})}-1<\int\gamma_{h}^{*}\lambda\,.$ (**)
This can easily be arranged as follows. As the quantity at the right hand side only depends on the thickness of the Weinstein handle realizing the contact surgery (which has been fixed), up to homotopy, supported in an open set $U\subset V$ so that $U\times S^{1}$ avoids the contact surgery, of $f$ among Morse functions satisfying all the properties listed above, we can move all the critical points of $f$ but the minimum to be so near to the boundary of $V$, where $f=1$, so that the desired inequality is satisfied. Note that such an open set $U$ and homotopy exist by the assumption done before that the contact surgery is done in the region $f^{-1}((a,b))\times S^{1}$, where $a$ is the global minimum of $f$ and $b$ is the image of the index $1$ critical points (recall again that $f$ is self-indexing).
@@
Let’s now start to arrange the desired almost complex structure in order to ensure that we can indeed use the energy bound (*4) described before. On the neighborhood $\partial V\times\mathbb{D}^{2}$ of the binding, the contact structure is given by
$\xi_{\partial V}\oplus\langle\,\partial_{x}+yR_{\lambda_{V}},\partial_{y}-xR_{\lambda_{V}}\,\rangle,$
@@
where $\xi_{\partial V}\subset T\partial V$ is the contact structure $\ker\lambda_{V}$, $R_{\lambda_{V}}$ is the Reeb vector field of $(\partial V,\lambda_{V})$ and $x,y$ are coordinates on $\mathbb{D}^{2}$. Moreover, the Reeb vector field $R$ of $\alpha$ on the neighborhood $\partial V\times\mathbb{D}^{2}$ of the binding is just $R_{\lambda_{V}}$. We then use an almost complex structure $J$ on $\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}$ satisfying the following properties:
@@
$J:\xi_{\partial V}\to\xi_{\partial V}\text{ is compatible with }\mathrm{d}\lambda_{V},\quad J(\partial_{x}+yR_{\lambda_{V}})=\partial_{y}-xR_{\lambda_{V}}\,\text{ and }\,J(R_{\lambda_{V}})=-\partial_{t}\,.$
@@
Note that with this choice of $J$, the natural projection $\pi_{D}$ from $\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2}$ to $\mathbb{D}^{2}$ is $(J,i)$-holomorphic, where $i$ is the standard complex structure on $\mathbb{D}^{2}$. In particular, we can indeed use the energy lower bound described in (*4).
Let’s now describe the Hamiltonian used. First on the filling $W_{0}=V\times\mathbb{D}^{2}$ of $\partial(V\times\mathbb{D}^{2})$ (i.e. before connect sum with $M$), one can consider a Hamiltonian $H$ with slope $2$ which vanishes on the filling. There are then two non-constant Hamiltonian orbits $\check{\gamma}_{p},\hat{\gamma}_{p}$ from each one of the previously described Reeb orbit $\gamma_{p}$’s with $p$ critical point of the Morse function $f$. Then, by [Zho22a, Proposition 2.6], for every
$\beta\in\operatorname{image}(H^{n}(V\times\mathbb{D})\to H^{n}(\partial(V\times\mathbb{D})))\simeq H^{n}(V)\,,$
there exists $\sum a_{i}\check{\gamma}_{p_{i}}$ in $SH^{*}_{+}$ and $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=\beta$ with $\mathrm{ind}(p_{i})=n$; in fact, one simply has $\beta=\sum a_{i}p_{i}$.
We now look at the connected sum with $M$ described before, and use the special contact form $\lambda$ on $M\#\partial(V\times\mathbb{D}^{2})$ which has been previously described. Recall in particular that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period $<2$ are just the $\gamma_{p}$’s and (some of) the multiple covers of the simple Reeb orbit $\gamma_{h}$ in the belt sphere of the handle, and that the inequality (**4) holds for any $p_{i}$ which is not the minimum point of $f$.
Note that, since the $p_{i}$’s in the above identity $\beta=\sum a_{i}p_{i}$ all have index $n$, the associated $\gamma_{p_{i}}$’s have the minimum period among all the Reeb orbits of the form $\gamma_{p}$. In particular, (2), (3) and the first claim of (4) will directly follow from action considerations, provided we can rule out the possibility of the $\gamma_{h}^{k}$’s, or the corresponding Hamiltonian orbits, appearing as negative asymptotics. To do this, we use an argument inspired from [CDvK16, Lemma 5.5], which leverages the energy lower bound described in (*4) in the binding region. (In fact, our situation is simpler than that in [CDvK16], as we are working with the trivial open book $\partial(V\times\mathbb{D})$.)
As the only Reeb orbits of action $<D$ on the connect sum have been arranged to be the ones from $\partial(V\times D^{2})$ and the iterates of the Reeb orbit on the belt sphere of the handle, for the moduli spaces in the conclusion of the proposition, it is enough to look at $\overline{\gamma}_{h}^{k}\in\{\check{\gamma}_{h}^{k},\hat{\gamma}_{h}^{k}\}$ and $\Gamma$ be a multiset of Reeb orbits taken from the $\gamma_{h}^{k}$’s. Then, for any curve $u$ in ¯$\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h}^{k},\Gamma)$, we have
@@
$\int_{S^{1}}\check{\gamma}_{p_{i}}^{*}\lambda-\int_{S^{1}}(\gamma_{h}^{k})^{*}\lambda-\sum_{\gamma\in\Gamma}\int_{S^{1}}\gamma^{*}\lambda\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}u^{*}\mathrm{d}\lambda\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}(\pi_{\mathbb{D}}u)^{*}\mathrm{d}(\frac{r^{2}}{2\pi}\mathrm{d}\theta)=1\,,$
@@
where the second inequality follows from (*4), and the last equality follows from the fact the fact that the linking number of the $\gamma_{p_{i}}$’s and of the $\gamma_{h}^{k}$’s around the binding $\partial V$ are $1$ and 00 respectively. Hence we arrive at a contradiction with the inequality (**4), provided such a curve $u$ in ¯$\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h}^{k},\Gamma)$ exists. This establishes (2), (3) and the first claim of (4).
@@
Lastly, using the Viterbo transfer map we see that, in the connected sum, $\sum a_{i}\check{\gamma}_{p_{i}}$ on $M\#\partial(V\times\mathbb{D})$ is mapped to $\sum a_{i}\check{\gamma}_{p_{i}}$ on $\partial(V\times\mathbb{D})$. Hence $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=0\oplus\beta$ for the Liouville filling $F\natural(V\times\mathbb{D})$. Combining with the just proved emptiness of the moduli spaces of the points (2), (3) and (4) in the statement, we then get the second claim of (4) by a neck-stretching argument, as done in Case 1. ∎
Proof of Theorem 4.1.
As in the statement, assume $W$ is a strong filling of $Y$. By the universal coefficient theorem, it is equivalent to prove the dual statement that $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$ is surjective onto $0\oplus\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$. By [CE20, Laz20b], there is a Weinstein cobordism from $M$ to $M^{\prime}$, such that $M^{\prime}$ is Weinstein fillable. Combining with [BCS15, Proposition 3.4], we can assume the Weinstein cobordism has vanishing first Chern class if $M$ does and $\dim M=5$. Then we have a Weinstein cobordism from $Y$ to $Y^{\prime}:=M^{\prime}\#\partial W_{0}$, yielding a strong filling $W^{\prime}$ of $Y^{\prime}$. It is clear that the claim follows from the surjective property for $W^{\prime},Y^{\prime}$. In other words, it suffices to prove for Liouville fillable $M$. For this purpose, let $\beta\in\operatorname{image}(H^{n}(W_{0};\mathbb{Q})\to H^{n}(\partial W_{0};\mathbb{Q}))$; we need to prove that $0\oplus\beta$ is in the image of $H^{n}(W;\mathbb{Q})\to H^{n}(Y;\mathbb{Q})$.
We will apply Proposition 4.3; let in particular $H$ and $\sum a_{i}x_{i}$ be the resulting Hamiltonian and $\mathbb{Q}$-linear combination of non-constant periodic orbits. Viewing $H$ as a Hamiltonian on $\widehat{W}$, then by applying neck-stretching along the contact boundary, we see that, for a sufficiently stretched almost complex structure, $\sum a_{i}x_{i}$ represents a class in $SH_{+}^{*}(W)$ (with Novikov field coefficient over $\mathbb{Q}$) by (2) and (3) of Proposition 4.3. Moreover, (4) of Proposition 4.3 implies that $\sum a_{i}x_{i}$ is mapped to $0\oplus\beta$ through $\delta_{\partial}\colon SH^{*}_{+}(W)\to H^{*+1}(W)\to H^{*+1}(Y)$. In particular, $0\oplus\beta$ lies in the image of $H^{*+1}(W)\to H^{*+1}(Y)$, as desired. ∎

### 14 Mar 2026 v5

@@ 4.7. The subcritical case
    The set of all Reeb orbits with period <2<2 is just $\{\gamma_{p}\;|\;p\in\mathrm{Crit}(f)\}$, see [Zho23b, Proposition 2.2].
We next explain how to construct an admissible almost complex structure to ensure the energy bound in (**3) in the following lemma, which we will then use in the proof of Proposition 4.5 in the subcritical case.
Lemma 4.6.
Let $V$ be a $2n$-dimensional Weinstein domain for $n\geq 2$. Consider also $W_{0}=V\times\mathbb{D}^{2}$ with its natural product symplectic structure and $(M,\xi)$ any contact manifold.
Then, there are
  • a Morse function $f$ on $V$ as in the definition of the contact form on $\partial(V\times\mathbb{D}^{2})$ above with maximal Morse index $n$,
  • a contact form $\alpha$ on $Y=\partial W_{0}\#M$,
  • an admissible almost complex structure $J$ on the symplectization $\mathbb{R}_{s}\times Y$
satisfying the following properties:
@@
    the Reeb orbits of $\alpha$ that have period <2<2 are
    • either from the set $\{\gamma_{p}\mid p\in\mathrm{Crit}(f)\}$, where each $\gamma_{p}$ has period $1/f(p)$ and has linking number $1$ in $Y$ with the binding $\partial V$ of the natural trivial open book on $\partial(V\times\mathbb{D}^{2})$,
    • or a multiple cover of a simple Reeb orbit $\gamma_{h,i}$ in the belt sphere of the $1$-handle in the connected sum, which has zero linking number with the binding;
  2. (2)
    the simple Reeb orbit $\gamma_{h,i}$ satisfies the action constraint
    $\int\gamma_{h,i}^{*}\alpha>\frac{1}{f(p)}-1\,$, (*)
    for all critical points $p$ of $f$ that are not the minimum and $1\leq i\leq n$;
  3. (3)
    all the maps $u$ defined on a subset of $\mathbb{R}\times\mathbb{S}^{1}$ that are contained in $\mathbb{R}_{s}\times\partial V\times\mathbb{D}^{2}$ and satisfy the Floer equation with respect to a Hamiltonian depending only on $s$, project via the natural projection $\pi_{\mathbb{D}}\colon\mathbb{R}_{s}\times\partial V\times\mathbb{D}^{2}\rightarrow\mathbb{D}^{2}$ to a holomorphic curve (with respect to the standard $i$ on $\mathbb{D}^{2}$), and we have
    $\int u^{*}{\mathrm{d}}\alpha\geq\int(\pi_{\mathbb{D}}u)^{*}{\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)\geq 0.$ (**)
Recall that, according to [Cie02b], any subcritical Weinstein domain $W_{0}$ is Liouville homotopic to the symplectic product $V\times\mathbb{D}^{2}$ for some Weinstein manifold $V$. In particular, the assumptions in Lemma 4.6 do indeed correspond to the hypothesis of Theorem 4.1 (in the subcritical $W_{0}$ case).
We now give a proof of this preparatory lemma.
Proof of Lemma 4.6.
The connected sum $M\#\partial(V\times\mathbb{D}^{2})$ is given by contact surgery of index $1$ on the disjoint union $M\sqcup\partial(V\times\mathbb{D}^{2})$; equivalently, it is the convex contact boundary of a Weinstein $1$-handle attachment on the disjoint union of $M$ with $\partial(V\times\mathbb{D}^{2})$. This surgery (or handle attachment) can be performed inside the subset $V\times\mathbb{S}^{1}\subset\partial(V\times\mathbb{D}^{2})$, away from every critical point of $f$ times the $\mathbb{S}^{1}$ factor. We first arrange property (2). For this, we start with a function $f_{0}$ on $V$ satisfying all the properties (of the function that was then denoted $f$) described right before the statement of this lemma. We now explain in steps how such functions can be modified in order to arrange property (2).
  1. (1)
    Let $x_{0}$ be the unique minimum point of $f_{0}$. We may assume the Liouville form on $\lambda_{V}$ on a ball $B$ around $x_{0}$ is the standard radical Liouville form on $\mathbb{C}^{n}$. Then $(V\backslash B,\lambda_{V})$ is a Liouville cobordism from the standard contact sphere with the standard contact form.
  2. (2)
    We may assume the $f_{0}$ near $\partial B$ only depends on the radius. More precisely, those can be achieved by first finding a chart near $x_{0}$ with $f_{0}$ depending only on the radius, then changing the Liouville form $\lambda_{V}$ to the standard one via a compactly supported Liouville homotopy with the same symplectic form.
  3. (3)
    We attach the handle to $(B\times\mathbb{S}^{1},\lambda_{V}+\frac{1}{2\pi f_{0}}{\mathrm{d}}\theta)$. In particular, the size of the handle is fixed at this stage.
  4. (4)
[…]
  5. (5)
[…]
  6. (6)
    We need to argue that $f$ on $V^{\prime}$ has the same property discussed before the lemma. The most important property is regarding the short Reeb orbits. By the discussion in [Zho23b, §2.1] and [Zho21b, Proposition 6.4], this essentially requires the Hamiltonian vector field of $1/f$ w.r.t. the symplectic form ${\mathrm{d}}\lambda_{V}$ 666[Zho21b, Proposition 6.4] uses ${\mathrm{d}}(\lambda_{V}/f)$ and a horizontal lift. In our case here, it is straightforward to compute that the formula for Reeb vector in [Zho21b, Proposition 6.4] is $(1/f-\lambda_{V}(X_{1/f}))^{-1}(\partial_{\theta}-X_{1/f})$, where $X_{1/f}$ is tha Hamiltonian vector field of $1/f$ using ${\mathrm{d}}\lambda_{V}$. to have no non-constant orbits with short periods. This is why we need the $C^{1}$-smallness of ${\mathrm{d}}f$ outside of a collar neighborhood of $\partial V$. The procedure of inserting a large trivial cobordism, i.e. blowing up the symplectic form, will scale down the Hamiltonian vector field even further. Therefore, it will not break the property on short Reeb orbits. In the following, we will return to the notation of $(V,\lambda_{V})$, which is considered as having gone through the modification above.
By scaling up the contact form on $M$ and Lemma 4.3, we can arrange that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period <2<2 are just the $\gamma_{p}$’s and (some of) the multiple covers of simple Reeb orbits $\gamma_{h,i}$ on the belt sphere of the handle, thus proving property (1) modulo the claim on linking numbers. As the binding $V\times\{0\}$ is null homologous, the linking number can be defined using any bounding disk of the Reeb orbit. For $\gamma_{p}$, we can use the natural disk from the $\mathbb{D}^{2}$ factor, and for $\gamma_{h,i}^{k}$, we can use a bounding disk contained in the handle region to deduce the claim concerning the linking number.
@@
All that is left to do is describe the desired almost complex structure $J$, and prove property (3). On the neighborhood $(\partial V\times\mathbb{D}^{2},\alpha=\lambda_{V}+\frac{r^{2}}{2\pi}{\mathrm{d}}\theta)$ of the binding, the contact structure is given by
$\xi_{\partial V}\oplus\langle\,\partial_{x}+\frac{y}{2\pi}R_{\lambda_{V}},\partial_{y}-\frac{x}{2\pi}R_{\lambda_{V}}\,\rangle$,
@@
where $\xi_{\partial V}\subset T\partial V$ is the contact structure $\ker\lambda_{V}$, $R_{\lambda_{V}}$ is the Reeb vector field of $(\partial V,\lambda_{V})$ and $x,y$ are coordinates on $\mathbb{D}^{2}$. Moreover, the Reeb vector field $R$ of $\alpha_{0}$ on the neighborhood $\partial V\times\mathbb{D}^{2}$ of the binding is just $R_{\lambda_{V}}$. We then use an almost complex structure $J$ on $\mathbb{R}_{s}\times\partial V\times\mathbb{D}^{2}$ satisfying the following properties:
@@
$J:\xi_{\partial V}\to\xi_{\partial V}\text{ is compatible with }\mathrm{d}\lambda_{V},\quad J(\partial_{x}+\frac{y}{2\pi}R_{\lambda_{V}})=\partial_{y}-\frac{x}{2\pi}R_{\lambda_{V}}\,\text{ and }\,J(R_{\lambda_{V}})=-\partial_{s}\,.$
@@
Note in particular that $\pi_{\mathbb{D}}$ is $(J,i)$ holomorphic, where $i$ is the standard complex structure on $\mathbb{D}^{2}$.
Moreover, if we have a Hamiltonian function $H$ on $\mathbb{R}_{s}\times\partial V\times\mathbb{D}^{2}$ that only depends on the $s$-coordinate, its Hamiltonian vector field is parallel to $R_{\lambda_{V}}$. If $u$ solves the Floer equation in $\mathbb{R}_{s}\times\partial V\times\mathbb{D}^{2}$, namely
$\partial_{s}u+J(\partial_{t}u-X_{H})=0\,$,
then $\pi_{\mathbb{D}}u$ is a holomorphic map to $\mathbb{D}$. Note that
@@
$u^{*}{\mathrm{d}}\alpha(\partial_{s}u,\partial_{t}u)={\mathrm{d}}\alpha(\partial_{s}u,J\partial_{s}u)={\mathrm{d}}\lambda_{V}(\pi_{\xi_{\partial V}}\partial_{s}u,\pi_{\xi_{\partial V}}J\partial_{s}u)+{\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\pi_{\mathbb{D}}\partial_{s}u,\pi_{\mathbb{D}}J\partial_{s}u)$
where $\pi_{\xi_{\partial_{V}}}$ is the natural projection to $\xi_{\partial V}$ and $\pi_{\mathbb{D}}$, by a bit abuse of notation, also denotes the natural projection to the $\mathbb{D}^{2}$-direction on the tangent bundle. Now, we can compute
${\mathrm{d}}\lambda_{V}(\pi_{\xi_{\partial V}}\partial_{s}u,\pi_{\xi_{\partial V}}J\partial_{s}u)={\mathrm{d}}\lambda_{V}(\pi_{\xi_{\partial V}}\partial_{s}u,J\pi_{\xi_{\partial V}}\partial_{s}u)\geq 0$
and
${\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\pi_{\mathbb{D}}\partial_{s}u,\pi_{\mathbb{D}}J\partial_{s}u)$ $=$ ${\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\pi_{\mathbb{D}}\partial_{s}u,i\pi_{\mathbb{D}}\partial_{s}u)$
$=$ ${\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\partial_{s}(\pi_{\mathbb{D}}\circ u),i\partial_{s}(\pi_{\mathbb{D}}\circ u))$
$=$ ${\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\partial_{s}(\pi_{\mathbb{D}}\circ u),\partial_{t}(\pi_{\mathbb{D}}\circ u))$
$=$ $(\pi_{\mathbb{D}}\circ u)^{*}{\mathrm{d}}\left(\frac{r^{2}}{2\pi}{\mathrm{d}}\theta\right)(\partial_{s}u,\partial_{t}u)\,.$
Therefore, (**3) holds. This concludes the proof of property (3), and hence of Lemma 4.6. ∎
Proof of Proposition 4.5 for $W_{0}$ subcritical.
The contact form $\alpha$ (depending on a function $f$ on $V$ as above) and almost complex structure $J$ are exactly those given by Lemma 4.6; we also use here $\gamma_{p}$ and $\gamma_{h,i}$ as given by said lemma. We also fix a Morse function $g$ on $\partial(V\times\mathbb{D}^{2})$. We now proceed to describe the desired Hamiltonian, $\mathbb{Q}$-linear combination, and then to prove the desired properties.
We start by describing the Hamiltonian. First, on the filling $W_{0}=V\times\mathbb{D}^{2}$ of $\partial(V\times\mathbb{D}^{2})$, one can consider a Hamiltonian $H$ which is linear and with slope $2$ outside a cylindrical neighborhood of the boundary, which vanishes on a slightly shrinked version of the filling, and which interpolates between these two behaviors along the cylindrical end, where it depends only on the Liouville coordinate. To utilise the positivity of contact energy in (10), we will use a Hamiltonian that only depends on $r$. This requires the cascades model to define the symplectic cohomology to deal with Morse-Bott families from the $S^{1}$-symmetry following [BO09b] or see [Zho23b, §2.2]. From each one of the Reeb orbits $\gamma_{p}$’s from Lemma 4.6, with $p$ a critical point of the Morse function $f$, we associate with two generators $\check{\gamma}_{p},\hat{\gamma}_{p}$ for symplectic cohomology. The moduli spaces in the cascades model, analogous to those in §4.4 were described in [Zho23b, §2.5]. In practice, one can neglect the difference between cascades and non-cascades versions of those in §4.4 as they enjoy the same properties (virtual dimensions, energy, etc.).
As far as the desired $\mathbb{Q}$-linear combination is concerned, we also point out that, for every
$\beta\in\operatorname{image}(H^{n}(V\times\mathbb{D})\to H^{n}(\partial(V\times\mathbb{D})))\cong H^{n}(V)\,$,
we can first represent $\beta=\sum a_{i}p_{i}$ in Morse cohomology, with $\mathrm{ind}(p_{i})=n$. Then, according to [Zho23b, Proposition 2.6], $\sum a_{i}\check{\gamma}_{p_{i}}$ in $SH^{*}_{+}$ satisfies $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=\beta$.
We now look at the connected sum of $\partial W_{0}$ with $M$ described before, and use the special contact form $\alpha$ from Lemma 4.6 as described above. Recall in particular that the Reeb orbits on $M\#\partial(V\times\mathbb{D})$ of period <2<2 are just the $\gamma_{p}$’s and (some of) the multiple covers of simple Reeb orbits $\gamma_{h,j}$ in the belt sphere of the handle, and that the inequality (*2) holds for any critical point $p_{i}$ which is not the minimum of $f$. Unlike the argument in the flexible case, we will use a Hamiltonian on the Liouville completion of $(V\times\mathbb{D}^{2})\natural F$ that has slope $2$ outside a cylindrical neighborhood of the boundary and vanishes on a slightly smaller copy of the filling.
Note that, since the $p_{i}$’s in the above identity $\beta=\sum a_{i}p_{i}$ all have index $n$, the associated $\gamma_{p_{i}}$’s have the minimum period among all the Reeb orbits of the form $\gamma_{p}$. In particular, (1), (2), and the first claim of (3) will directly follow from action considerations, provided we can rule out the possibility of the $\gamma_{h,j}^{k}$’s, or the corresponding Hamiltonian orbits, appearing as negative asymptotics. To do this, we use an argument inspired by [CDvK16, Lemma 5.5], which leverages the energy lower bound described in (**3) of Lemma 4.6 in the binding region. (In fact, our situation is simpler than that in [CDvK16], as we are working with the trivial open book $\partial(V\times\mathbb{D})$.)
As the only Reeb orbits of action <2<2 on the connect sum have been arranged to be the ones from $\partial(V\times\mathbb{D}^{2})$ and the iterates of the Reeb orbit on the belt sphere of the handle, for the moduli spaces in the conclusion of the proposition, it is enough to look at $y=\overline{\gamma}_{h,j}^{k}\in\{\check{\gamma}_{h,j}^{k},\hat{\gamma}_{h,j}^{k}\}$ and $\Gamma$ a multiset of Reeb orbits taken from the $\{\gamma_{h,j}^{k}\}_{k,j}$. Then, for any curve $u$ in $\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h,j}^{k},\Gamma)$, we have
$\int_{S^{1}}\check{\gamma}_{p_{i}}^{*}\alpha-\int_{S^{1}}(\gamma_{h,j}^{k})^{*}\alpha-\sum_{\gamma\in\Gamma}\int_{S^{1}}\gamma^{*}\alpha\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}u^{*}\mathrm{d}\alpha\;\geq\;\int_{u^{-1}(\mathbb{R}_{t}\times\partial V\times\mathbb{D}^{2})}(\pi_{\mathbb{D}}u)^{*}\mathrm{d}\left(\frac{r^{2}}{2\pi}\mathrm{d}\theta\right)=1\,$, (10)
@@
where the first inequality follows from Stokes’ theorem as $u^{*}{\mathrm{d}}\alpha\geq 0$ on the whole curve as $H$ only depends on $r$ in this region, the second inequality from (**3), and the last equality from the fact that the linking number of the $\gamma_{p_{i}}$’s and of the $\gamma_{h,j}^{k}$’s around the binding $\partial V$ are $1$ and 0 respectively. Hence we arrive at a contradiction with the inequality (*2), provided such a curve $u$ in $\overline{{\mathcal{M}}}_{Y,H}(\check{\gamma}_{p_{i}},\overline{\gamma}_{h,j}^{k},\Gamma)$ exists. This establishes (1), (2) and the first claim of (3).
@@
Lastly, using the Viterbo transfer map we see that, in the connected sum, $\sum a_{i}\check{\gamma}_{p_{i}}$ on $M\#\partial(V\times\mathbb{D})$ is mapped to $\sum a_{i}\check{\gamma}_{p_{i}}$ on $\partial(V\times\mathbb{D})$. This can be seen using the action filtration as $\gamma_{p_{i}}$ has the minimal period from those $\gamma_{p}$, the transfer map from $\check{\gamma}_{p_{i}}$ are built from reparameterization of the trivial cylinder, as explained in [Zho23a, §4.2, (4.3)]. Hence $\delta_{\partial}(\sum a_{i}\check{\gamma}_{p_{i}})=0\oplus\beta$ for the Liouville filling $F\natural(V\times\mathbb{D})$. Combining with the just proved emptiness of the moduli spaces of the points (1), (2), and (3) in the statement, we then get the second claim of (3) by a neck-stretching argument by applying the proof of Proposition 4.5 in the previous case mutatis mutandis. ∎